### v1

@@ Abstract
We study real-valued mean-field backward stochastic differential equations (BSDEs) of the form
$Y_{t}=\xi+\int_{t}^{T}\widetilde{\mathbb{E}}\,g(s,Z_{s},\tilde{Z}_{s})\,ds-\int_{t}^{T}Z_{s}\,dW_{s},$
where $\tilde{Z}$ denotes an independent copy of $Z$ and $\widetilde{\mathbb{E}}$ the expectation with respect to $\tilde{Z}$. Under a separately quadratic growth assumption (Hq) on the generator $g$ in (Z,Z~)($\tilde{Z}$), together with a bounded terminal condition, we prove existence and uniqueness of solutions in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates: we first obtain uniform $\mathbb{S}^{\infty}$-bounds for $Z$ in a Lipschitz setting with bounded Malliavin derivative, and then pass to the quadratic case by approximation, using a stability result in $\mathbb{S}^{2}\times\mathbb{H}^{2}$. This closes the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean.
In the second part of the paper, we extend the framework to generators of the form $g(t,z,\tilde{z},\bar{z})$ satisfying a fully coupled quadratic condition (Hq). In this general regime we establish existence and uniqueness under a smallness condition on the centered terminal variable. As an application, we solve a mean-field exponential utility maximization problem with a collective liability, thereby generalizing the classical utility maximization framework of hu2005utility to a fully coupled quadratic mean-field setting.
@@ 1 Introduction
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}dW_{s},$ (1.1)
@@
The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate. The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$. The form of the equation is as follows:
$Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}dW_{s},\quad t\in[0,T],$ (1.2)
@@
under the conditions that the generator $F$ is Lipschitz continuous and $\xi$ is square integrable. For a comprehensive presentation of the theory, see the review by el1997backward books by CohenStocal2015 or ZhangBSDEbook2017. Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the coefficients or the conditions on the terminal value $\xi$. Consequently, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations.
@@
BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention. These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems with exponential utility functions hu2005utility or in risk minimization problems for the entropic risk measure. Kobylanski:2000cy first provided the existence and uniqueness result for scalar-valued quadratic BSDEs with bounded terminal values, where $F$ is quadratic with respect to its $Z$ component. A detailed review of the literature can be found in el2008bsde. briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods. Investigations by briand2006bsde extend the previous results to unbounded unbounded terminal conditions. Some studies generalized to the multidimensional case by hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
@@
Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles. These equations date back to kac1956foundations’s work on the Boltzmann and Vlasov equations in the 1950s. Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame. Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$-particle systems, and investigated them using probabilistic methods. This has led to a growing interest in mean-field SDEs and BSDEs. For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions and a comparison theorem for mean-field BSDEs with Lipschitz continuous generators and square integrable terminal conditions. More recently, hao2022mean explored the solvability of local and global solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of local and global solutions for multidimensional BSDEs with generators depending on the expectation of both variables. Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
@@
pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus. In fact, Buckdahn2009MFBSDE states that, in a Clark–Ocone type formula, the martingale density component of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component (see Proposition 5.3). The close relationship between Malliavin calculus and BSDEs stems from the use of Malliavin calculus to prove the regularity of trajectories. For example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth as one of the applications.
@@
Our methods are inspired by the works of tevzadze2008solvability and briand2013simple. The former studied the existence and uniqueness of solutions for multidimensional quadratic BSDEs, utilizing BMO martingale theory. The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus. Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean. Specifically, hao2022mean established local existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution. Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions set forth in hao2022mean, thereby proving the existence of a solution for quadratic mean-field BSDEs.
Our main contributions are as follows. First, we establish the existence and uniqueness of a solution when the generator exhibits separately quadratic growth in both $Z$ and the mean-field term $\tilde{Z}$ with bounded terminal condition $\xi$. Our approach builds upon the theory of BMO martingales and Malliavin calculus to effectively address the analytical challenges arising from the quadratic structure. A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature. Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2023MFBSDE and tevzadze2008solvability.
@@ 2.1 Preliminaries
By $\left|\cdot\right|$, we denote the Euclidean norm on $\mathbb{R}^{d}$, and for $p>1$, we denote by
  • $L^{p}$: all real-valued random variables $X$ s.t. $\|X\|_{L^{p}}^{p}:={\mathbb{E}}^{\mathbb{P}}\big[|X|^{p}\big]<\infty$,
@@
    $\left\|Y\right\|_{\mathbb{S}^{p}}:=\mathbb{E}\left[\sup_{0\leq r\leq T}\left|Y_{r}\right|^{p}\right]^{\frac{1}{p}}<\infty,$
@@
    $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{\omega\in\Omega}\sup_{0\leq t\leq T}\left|Y_{t}(\omega)\right|<\infty,$
  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\mathbb{E}\left[\left(\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right)^{\frac{p}{2}}\right]^{\frac{1}{p}}<\infty,$
  • $\mathbb{H}^{2}_{BMO}$: the subsets of $\mathbb{H}^{2}$ with predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\int_{0}^{\cdot}Z_{s}\cdot\,dW_{s}$ is a Bounded in Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{s}\right|^{2}\,ds|\mathcal{F}_{\tau}\right]\leq C^{2}.$
    The $\left\|\int_{0}^{\cdot}Z_{s}\cdot\,dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
  • Let $C$ be a closed subset of $\mathbb{R}^{n}$ and $a\in\mathbb{R}^{n}$. The distance between $a$ and the set $C$ is given by
@@
When a different underlying measure is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$. Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
@@
For $K>0$. There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds. Let $N$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
@@ 2.2 Statement of the main theorem
In this paper we consider the following general mean-field BSDE
@@ 2. $(\rm H_{q})$}
In the statement and proof of our main theorem we may, without loss of generality, assume that
@@
[…]
@@ 3 Terminal conditions with bounded Malliavin derivative
This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.5) under the assumption that the driver $g$ is Lipschitz continuous. These bounds become particularly useful when the terminal condition $\xi$ is Malliavin differentiable. Although this may appear to be a stronger condition, it serves a crucial role in our analysis. In the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques. Therefore, we initially work under stronger regularity assumptions to facilitate this reduction. For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition.
@@
    There exist constants $L_{z}>0$ such that:
@@
Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$. Therefore, the Novikov condition is satisfied, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
@@
Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.7. ∎

### v2

@@ Abstract
In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generator of quadratic growth in the control variable and the mean-field term. Under this assumption, together with a bounded terminal condition, we establish existence and uniqueness of solutions. Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates. The result bridges the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean. Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to generators satisfying a weaker quadratic condition on the generator. This relaxation is designed to accommodate the additional mean-field terms that arise in our utility maximization setting and that fall outside the scope of previous quadratic assumptions. Within this more general regime, we establish existence and uniqueness of solutions under a smallness condition on the terminal random variable. We then apply this extended theory to solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility to a fully coupled quadratic mean-field setting.
@@ 1 Introduction
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},$ (1.1)
@@
The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate. The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$ for
$Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],$ (1.2)
@@
under the conditions that the generator $F$ is Lipschitz continuous and the terminal condition $\xi$ is square integrable. Since then, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations. For a comprehensive presentation of the theory, see el1997backward, CohenStocal2015 or ZhangBSDEbook2017. Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the generator $F$ or the conditions on the terminal condition $\xi$.
@@
BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention. These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems in an incomplete market hu2005utility. Kobylanski:2000cy provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component. briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods. Investigations by briand2006bsde extend the previous results to unbounded terminal conditions. Some studies the multidimensional cases: see hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
@@
Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles. These equations date back to kac1956foundations’s work on the Boltzmann and Vlasov equations in the 1950s. Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame. Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods. This has led to a growing interest in mean-field BSDEs and FBSDEs. For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions. More recently, hao2022mean explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables. Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
@@
pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus. In fact, el1997backward and Buckdahn2009MFBSDE state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3). The result has been extending our understanding of quadratic BSDEs: for example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth.
@@
Our methods are inspired by the works of tevzadze2008solvability and briand2013simple. The former studied the existence and uniqueness of solutions for multidimensional quadratic BSDEs, utilizing BMO martingale theory. The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus. Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean. Specifically, hao2022mean established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution. Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in hao2022mean, thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the corresponding mean-field component $\tilde{Z}$.
Our main contributions are as follows. First, we establish the existence and uniqueness of a solution when the generator exhibits separately quadratic growth in both $Z$ and the mean-field term $\tilde{Z}$ with bounded terminal condition $\xi$. Our approach builds upon the theory of BMO martingales and Malliavin calculus to effectively address the analytical challenges arising from the quadratic structure. A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature. Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2024MFBSDE and tevzadze2008solvability.
@@ 2.1 Preliminaries
  • $L^{p}$: all real-valued random variables $X$ s.t. $\|X\|_{L^{p}}:=\left({\mathbb{E}}^{\mathbb{P}}\big[|X|^{p}\big]\right)^{\frac{1}{p}}<\infty$,
@@
    $\left\|Y\right\|_{\mathbb{S}^{p}}:=\left(\mathbb{E}\left[\sup_{0\leq r\leq T}\left|Y_{r}\right|^{p}\right]\right)^{\frac{1}{p}}<\infty,$
@@
    $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{0\leq r\leq T}\left\|Y_{r}\right\|_{L^{\infty}}=\sup_{\omega\in\Omega}\sup_{0\leq r\leq T}\left|Y_{r}(\omega)\right|<\infty,$
  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
  • $\mathbb{H}^{2}_{BMO}$: a subset of $\mathbb{H}^{2}$ with $Z$ satisfying $Z\cdot W:=\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ is a Bounded Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\Big|\mathcal{F}_{\tau}\right]\leq C^{2}.$
    The $\left\|Z\right\|_{\mathbb{H}^{2}_{BMO}}:=\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
  • For a closed subset $C$ of $\mathbb{R}^{n}$ and $a\in\mathbb{R}^{n}$, the distance between $a$ and the set $C$ is defined by
@@
When a different underlying measure other than $\mathbb{P}$ is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$. Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
@@
For $K>0$. There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds. Let $N$ be a one-dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
@@
Definition 2.3 (sliceability).
Let $M$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ on $[0,T]$. A random partition of $[0,T]$ is a finite sequence of stopping times $0=\tau_{0}\leq\tau_{1}\leq\cdots\leq\tau_{m}=T$. For $\delta>0$, the index of sliceability of $M$ is the function
$N_{M}:(0,\infty)\longrightarrow\mathbb{N}\cup\{\infty\}$
defined as follows: $N_{M}(\delta)$ is the smallest integer $m$ such that there exists a random partition $(t_{k})_{k=0}^{m}$ of $[0,T]$ with
$\big\|M^{(k)}\big\|_{\mathrm{BMO}}\leq\delta,\qquad 1\leq k\leq m,$ (2.5)
where the kk–th increment $M^{(k)}$ is the continuous martingale
$M^{(k)}_{t}:=M_{t\wedge\tau_{k}}-M_{t\wedge\tau_{k-1}},\qquad 0\leq t\leq T.$
If no such $m$ exists, we set $N_{M}(\delta)=\infty$. We say that $M$ is $\delta$–sliceable if $N_{M}(\delta)<\infty$, and simply sliceable if $M$ is $\delta$–sliceable for every $\delta>0$.
@@ 2.2 Statement of the main theorem
In this paper, we consider the following general mean-field BSDE
@@ 2. $(\rm H_{q})$}
Example 2.4.
The following $g$ satisfies the condition $(\mathrm{H}_{q})$:
[…]
where $l$ is Lipschitz in (z,z~)($\tilde{z}$), and $q_{1}$ and $q_{2}$ are quadratic polynomials in $z$ and $\tilde{z}$, respectively. Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as hao2022mean.
@@
In the statement and proof of our main theorem we may, without loss of generality, we can assume that
@@
solves the original BSDE. Thus there is a one-to-one correspondence in terms of existence and uniqueness. From this point on, we therefore work directly with the shifted generator, still denoted by $g$, and silently assume $g(t,0,\tilde{z})=0$ for all $(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}$.
@@ 3 Terminal conditions with bounded Malliavin derivative
This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.6) under the assumption that the driver $g$ is Lipschitz continuous and the terminal condition $\xi$ is Malliavin differentiable. These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques. Therefore, we initially work under stronger regularity assumptions to facilitate this reduction. For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition and condition on $\xi$ will be introduced in Propisiton 3.1 below.
@@
    There exist a constant $L_{z}>0$ such that:
@@
Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$. Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]$ is measurable to $mathcal{F_{s}}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
@@
Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.8. ∎